OpenStudy (anonymous):

A bag contains 3 black marbles, 1 yellow marble, and 6 blue marbles. Two marbles are drawn successively from the bag without replacement. What is the probability that both marbles are blue?

OpenStudy (anonymous):

\[\frac{6}{10}\times \frac{5}{9}\]